OpenStudy (anonymous):

Okay I want to know did I get this correct? Directions: Factor each polynomial The problem: m^2+7m-18 My Answer: m^2+9m-2m Directions:Factor each polynomial Problem: 2x^2-3x-5 My Answer: Couldnt get one

OpenStudy (cwrw238):

m^2+7m-18 - you need 2 numbers whose product is -18 and whose sum is + 7 the answer is of the form (x + a)(x + b) where a and b are the 2 numbers

OpenStudy (cwrw238):

*m instead of x

OpenStudy (cwrw238):

think of some factors of -18 which will give you +7 when you add them

OpenStudy (anonymous):

9 and -2?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

so the factors are (m + 9)(m - 2)

OpenStudy (anonymous):

Yeah I got that...so my answer just wasnt correct?

OpenStudy (cwrw238):

your original answer was wrong

OpenStudy (cwrw238):

second one can be factored also:- 2x^2-3x-5 = 2x^2 -5x + 2x - 5 = x(2x - 5) + 1(2x - 5) = (2x - 5)(x + 1)